there were 20 questions.5 marks were given for each correct answer and 3 marks will be deducted for each unanswered question or wrong answer. Anna only answered 6 questions correctly.Find Anna's score.

Answer by ikleyn(52847)

.
Positive marks: 6*5 = + 30.


Negative marks: (20-6)*(-3) = 14*(-3) = - 42.


The balance:    + 30 - 42 = -12.


Answer.  Anna' score is -12.